OpenStudy (anonymous):

What is the solution to the system of equations? Please show work. 3x+3y=10 -9x-9y=-30 Medal to first correct answer with proper work

OpenStudy (anonymous):

oh i know this one

OpenStudy (anonymous):

we can use the elimination method for these equations 3x+3y=10 <----multiply this eqation by 3 -9x-9y=-30 9x+8y=30 -9x-9y=-30 0=0 since both sides equal 0, the variables can be all real numbers

OpenStudy (anonymous):

sorry 9x+9y=30

OpenStudy (anonymous):

everyone, does that seem right? it seems right to me

OpenStudy (anonymous):

Its right

OpenStudy (anonymous):

thx :D

OpenStudy (anonymous):

lines are co-incident and has infinitely many soluions